A boy on a 2.0 kg skateboard initially at rest tosses an 8.0 kg jug of water in the forward direction. If the jug has a speed of 3.0 m/s relative to the ground and the boy and skateboard move in the opposite direction at 0.60 m/s, find the boy’s mass.


1
Expert's answer
2022-03-02T14:38:45-0500

According to the momentum conservation law, the magnitude of boy with skateboard momentum must be equal to the magnitude of jug's momentum:


"(m_b+m_s)v_1 = m_jv_2"

where "v_1 = 0.6m\/s, v_2 = 3m\/s,m_s = 2kg,m_j=8kg" and "m_b" is boy's mass. Thus. obtain:


"m_b = \dfrac{m_jv_2-m_sv_1}{v_1} = \dfrac{8\cdot 3 - 2\cdot 0.6}{0.6} = 38kg"

Answer. 38kg.